OpenStudy (anonymous):

HELP PLEASE!! What is the focus of the parabola given by −1/4(y+2)^2=(x−1)?

OpenStudy (anonymous):

a)(-1, 2) b)(1, -2) c)(0, -2) d)(0, 2)

OpenStudy (anonymous):

@Michele_Laino can you help please?

OpenStudy (michele_laino):

sorry is your parabola this? \[-\frac{ 1 }{ 4(y+2)^{2} }=x-1\]

OpenStudy (anonymous):

no, its the way i put it, in the question

OpenStudy (anonymous):

@adamanderson

OpenStudy (michele_laino):

is this? \[-\frac{ 1 }{ 4 }(y+2)^{2}=x-1\]

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

ok! Now you have to change your coordinate system, namely you have to perform a traslation, whose equation are: \[X=x-1,Y=y+2\] where X and Y are the new coordinates. In this new coordinate system, your parabola can be rewritten as below: \[X=-\frac{ 1 }{ 4 }Y ^{2}\]

OpenStudy (michele_laino):

coordinates of focus of this parabola are simply to find, and are: X=0, Y=0

OpenStudy (michele_laino):

but, please be careful, because we know X, and Y, and we have to know x and y instead. In order to find x, and y, we have to substitute X=0 Y=0, in the above transformation equations, namely: X=x-1, Y=y+2, so we have: \[0=x-1,0=y+2\] from which we can write: x=1, and y=-2 so b).

OpenStudy (anonymous):

took me awile but its (b

OpenStudy (anonymous):

thanks so much!!!! couldnt have done without you!! @Michele_Laino

OpenStudy (michele_laino):

thanks!!!!!